# Calculating the EFT likelihood via saddle-point expansion

### v1

@@ Appendix A The contribution of the saddle-point
$(S[\bar{\boldsymbol{\phi}}]-S[\bar{\boldsymbol{\phi}}_{g}])^{(3)}\supset{-{}}\int_{\boldsymbol{k}}\frac{P_{\epsilon_{g}\epsilon_{m}}(k)}{P_{\epsilon_{g}}(k)}\frac{\delta(\boldsymbol{k})\delta^{(2)}_{g,{\rm det}}[\delta](-\boldsymbol{k})}{P_{\rm L}(k)}\,\,.$ (A.9)
@@
×$\times\bigg{(}\frac{P_{\epsilon_{g}\epsilon_{m}}(p_{2})}{P_{\epsilon_{g}}(p_{2})}\delta(\boldsymbol{p}_{1})\big{(}\delta_{g}(\boldsymbol{p}_{2})-\delta_{g,{\rm det}}^{(1)}(\boldsymbol{p}_{2})\big{)}\bigg{)}(\boldsymbol{p}_{1}\to\boldsymbol{p}_{2})\bigg{]}\,\,.$ (A.10)
Appendix B The eigenvalues of $\mathcal{M}_{g}$
@@
$\begin{pmatrix}\boldsymbol{0}&\mathcal{M}^{*}_{g}\\ \mathcal{M}_{g}&\boldsymbol{0}\end{pmatrix}\begin{pmatrix}\chi_{n}(\boldsymbol{k})\\ {\chi}_{n}^{*}(\boldsymbol{k})\end{pmatrix}=m_{n}\begin{pmatrix}\chi_{n}(\boldsymbol{k})\\ {\chi}_{n}^{*}(\boldsymbol{k})\end{pmatrix}\,.$ (B.1)
@@
$P_{\epsilon_{g}}\equiv a\,,\,P_{\epsilon_{g}\epsilon_{m}}\equiv b\,,\,b_{1}D_{1}\equiv c\,,\,D_{1}\equiv d\,,\,P_{\rm in}^{-1}\equiv e\,.\,$ (B.2)
@@
$\begin{pmatrix}0&0&0&a&b&-ic\\ 0&0&0&b&0&-id\\ 0&0&0&-ic&-id&e\\ a&b&ic&0&0&0\\ b&0&id&0&0&0\\ ic&id&e&0&0&0\end{pmatrix}\begin{pmatrix}\chi_{n}(\boldsymbol{k})\\ {\chi}_{n}^{*}(\boldsymbol{k})\end{pmatrix}=m_{n}\begin{pmatrix}\chi_{n}(\boldsymbol{k})\\ {\chi}_{n}^{*}(\boldsymbol{k})\end{pmatrix}\,.$ (B.3)
@@
$m_{1}$ $=\frac{a-e}{3}-\frac{c_{1}}{3\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}-\frac{\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}{3}\,;$ (B.4)
$m_{2}$ $=\frac{a-e}{3}-\frac{c_{1}}{3\left(-\frac{1}{2}+\frac{\sqrt{3}i}{2}\right)\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}-\frac{(-\frac{1}{2}+\frac{\sqrt{3}i}{2})\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}{3}\,;$ (B.5)
$m_{3}$ $=\frac{a-e}{3}-\frac{c_{1}}{3\left(-\frac{1}{2}-\frac{\sqrt{3}i}{2}\right)\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3})}\right)^{\frac{1}{3}}}-\frac{(-\frac{1}{2}-\frac{\sqrt{3}i}{2})\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}{3}\,,$ (B.6)
@@
$c_{1}$ $=3ae+3b^{2}+3c^{3}+3d^{2}+(-a+e)^{3}\,,$ (B.7)
$c_{2}$ $=\frac{27ad^{2}}{2}-\frac{27b^{2}e}{2}-27bcd+\frac{9(e-a)(ae+b^{2}+c^{2}+d^{2})}{2}+(e-a)^{3}\,.$ (B.8)

### v2

@@ Appendix B The contribution of the saddle-point
$(S[\bar{\boldsymbol{\phi}}]-S[\bar{\boldsymbol{\phi}}_{g}])^{(3)}\supset{-{}}\int_{\boldsymbol{k}}\frac{P_{\epsilon_{g}\epsilon_{m}}(k)}{P_{\epsilon_{g}}(k)}\frac{\delta(\boldsymbol{k})\delta^{(2)}_{g,{\rm det}}[\delta](-\boldsymbol{k})}{P_{\rm L}(k)}\,\,.$ (B.9)
@@
×$\times\bigg{(}\frac{P_{\epsilon_{g}\epsilon_{m}}(p_{2})}{P_{\epsilon_{g}}(p_{2})}\delta(\boldsymbol{p}_{1})\big{(}\delta_{g}(\boldsymbol{p}_{2})-\delta_{g,{\rm det}}^{(1)}(\boldsymbol{p}_{2})\big{)}\bigg{)}(\boldsymbol{p}_{1}\to\boldsymbol{p}_{2})\bigg{]}\,\,.$ (B.10)
Appendix C The eigenvalues of $\mathcal{M}_{g}$
@@
$\begin{pmatrix}\boldsymbol{0}&\mathcal{M}^{*}_{g}\\ \mathcal{M}_{g}&\boldsymbol{0}\end{pmatrix}\begin{pmatrix}\chi_{n}(\boldsymbol{k})\\ {\chi}_{n}^{*}(\boldsymbol{k})\end{pmatrix}=m_{n}\begin{pmatrix}\chi_{n}(\boldsymbol{k})\\ {\chi}_{n}^{*}(\boldsymbol{k})\end{pmatrix}\,.$ (C.1)
@@
$P_{\epsilon_{g}}\equiv a\,,\,P_{\epsilon_{g}\epsilon_{m}}\equiv b\,,\,b_{1}D_{1}\equiv c\,,\,D_{1}\equiv d\,,\,P_{\rm in}^{-1}\equiv e\,.\,$ (C.2)
@@
$\begin{pmatrix}0&0&0&a&b&-ic\\ 0&0&0&b&0&-id\\ 0&0&0&-ic&-id&e\\ a&b&ic&0&0&0\\ b&0&id&0&0&0\\ ic&id&e&0&0&0\end{pmatrix}\begin{pmatrix}\chi_{n}(\boldsymbol{k})\\ {\chi}_{n}^{*}(\boldsymbol{k})\end{pmatrix}=m_{n}\begin{pmatrix}\chi_{n}(\boldsymbol{k})\\ {\chi}_{n}^{*}(\boldsymbol{k})\end{pmatrix}\,.$ (C.3)
@@
$m_{1}$ $=\frac{a-e}{3}-\frac{c_{1}}{3\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}-\frac{\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}{3}\,;$ (C.4)
$m_{2}$ $=\frac{a-e}{3}-\frac{c_{1}}{3\left(-\frac{1}{2}+\frac{\sqrt{3}i}{2}\right)\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}-\frac{(-\frac{1}{2}+\frac{\sqrt{3}i}{2})\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}{3}\,;$ (C.5)
$m_{3}$ $=\frac{a-e}{3}-\frac{c_{1}}{3\left(-\frac{1}{2}-\frac{\sqrt{3}i}{2}\right)\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3})}\right)^{\frac{1}{3}}}-\frac{(-\frac{1}{2}-\frac{\sqrt{3}i}{2})\left(\frac{1}{2}(c_{2}+\sqrt{4c_{2}^{2}-4c_{1}^{3}})\right)^{\frac{1}{3}}}{3}\,,$ (C.6)
@@
$c_{1}$ $=3ae+3b^{2}+3c^{3}+3d^{2}+(-a+e)^{3}\,,$ (C.7)
$c_{2}$ $=\frac{27ad^{2}}{2}-\frac{27b^{2}e}{2}-27bcd+\frac{9(e-a)(ae+b^{2}+c^{2}+d^{2})}{2}+(e-a)^{3}\,.$ (C.8)